OpenStudy (anonymous):

x^2 - 9/ x^2 + 5x + 6 * 1/ X+3 Anwsers: 1. 1 2. 1/x+2 3. x-3/(x+2)(x+3)

OpenStudy (campbell_st):

break the problem into parts \[x^2 - 9 = (x+3)(x-3) \] \[x^2 +5x +6 = (x +3)(x+2)\] I'll let you simplify the 1st part... then ,mutply your answer by \[1/(x +3)\] and you'll have the correct answer

OpenStudy (anonymous):

i cant figure that out?

OpenStudy (anonymous):

is it anwser 3?